OpenStudy (anonymous):

State whether the given measurements determine zero, one, or two triangles. A = 35°, a = 6, b = 13 Two One Zero

OpenStudy (anonymous):



OpenStudy (anonymous):

looks like there could be two, but we could check it

OpenStudy (anonymous):

\[\frac{\sin(35)}{6}=\frac{\sin(B)}{13}\] \[\sin(B)=\frac{13\sin(35)}{6}\] \[B=\sin^{-1}(\frac{13\sin(35)}{6})\] and now a calculator

OpenStudy (anonymous):

since \(\frac{13\sin(35)}{6}>1\) there is no way to take the inverse sine of this number therefore there is no such triangle

OpenStudy (anonymous):



OpenStudy (anonymous):

will it be zero

OpenStudy (anonymous):

yes there is no such triangle

OpenStudy (anonymous):

thanks for explaining the problem

OpenStudy (anonymous):

@satellite73 how did you know the height was 6?